OpenStudy (anonymous):

solve x^2 -23x/5 -2 = 0 (quadratic equation)

OpenStudy (anonymous):

\[x=\left( -b \pm \sqrt{b^2-4ac}\over2a \right)\]

OpenStudy (anonymous):

a=1, b=23/5, c=-2

OpenStudy (anonymous):

sorry. be = -23/5

OpenStudy (anonymous):

you should get x = 2/5 or x =-5